# OST-Bench: Evaluating the Capabilities of MLLMs in Online Spatio-temporal Scene Understanding

### v1

@@ Abstract
Recent advances in multimodal large language models (MLLMs) have shown remarkable capabilities in integrating vision and language for complex reasoning. While most existing benchmarks evaluate models under offline settings with a fixed set of pre-recorded inputs, we introduce OST-Bench, a benchmark designed to evaluate Online Spatio-Temporal understanding from the perspective of an agent actively exploring a scene. The “Online” aspect emphasizes the need to process and reason over incrementally acquired observations, while the “Spatio-Temporal” component requires integrating current visual inputs with historical memory to support dynamic spatial reasoning. OST-Bench better reflects the challenges of real-world embodied perception. Built on an efficient data collection pipeline, OST-Bench consists of 1.4k scenes and 10k question-answer pairs collected from ScanNet, Matterport3D, and ARKitScenes. We evaluate several leading MLLMs on OST-Bench and observe that they fall short on tasks requiring complex spatio-temporal reasoning. Under the online setting, their accuracy declines as the exploration horizon extends and the memory grows. Through further experimental analysis, we identify common error patterns across models and find that both complex clue-based spatial reasoning demands and long-term memory retrieval requirements significantly drop model performance along two separate axes, highlighting the core challenges that must be addressed to improve online embodied reasoning. To foster further research and development in the field, our codes, dataset, and benchmark are available at https://github.com/rbler1234/OST-Bench.
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs)bai2023qwentechnicalreport ; chen2024internvlscalingvisionfoundation ; liu2024improvedbaselinesvisualinstruction ; touvron2023llamaopenefficientfoundation ; lin2024vilapretrainingvisuallanguage ; zhang2024longcontexttransferlanguage ; li2024llavanextinterleavetacklingmultiimagevideo have shown remarkable capabilities in integrating vision and language for complex reasoning. However, most existing benchmarks scanrefer ; lyu2024mmscanmultimodal3dscene ; sqa3d ; scanqa ; sceneverse ; referit3d , evaluate models under offline settings, where reasoning is performed over a fixed set of pre-recorded inputs, such as reconstructed 3D scenes or images and videos, this does not capture the online nature of embodied tasks.
@@
To address this gap, we introduce OST-Bench, a benchmark designed to evaluate Online Spatio-Temporal understanding from the perspective of an agent actively exploring a scene. The term Online emphasizes the agent’s need to perceive, remember, and reason over incrementally received observations, rather than complete, pre-recorded scene data. The term Spatio-Temporal highlights the need to integrate current visual observation with historical memory to support dynamic spatial reasoning. To more accurately simulates real-world embodied perception, OST-Bench defines task from these perspectives: the agent, its surrounding environment, and their relationship, contains three main categories(Fig.2): (1) Agent State: the agent’s understanding of its own state, (2) Agent Visible Info: the agent’s dynamic interpretation of visible scene information, and (3)Agent-object Spatial Relationship: the agent’s dynamic understanding of spatial relationships with objects, all posed in an online, temporally grounded fashion. OST-Bench comprises 1.4k real-world scenes sourced from the test and validation splits of ScanNetscannet , Matterport3Dmp3d , and ARKitScenesarkitscenes , accompanied by 10k QA pairs covering a diverse range of subtypes. Our benchmark provides a rigorous testbed for assessing the online spatio-temporal reasoning ability of MLLMs in realistic, embodied settings.
@@
We evaluate leading MLLMs on OST-Bench and find its online spatio-temporal nature poses significant challenges for the models, even the most advanced models lag behind human performance by over 30%. Models perform poorly on tasks requiring complex spatio-temporal reasoning, with accuracy declining as exploration steps increase and memory grows under the online setting. Based on an in-depth experimental analysis, we observe a phenomenon which we term Spatio-temporal Reasoning Shortcut-when reasoning over long-term memory, models tend to avoid retrieving key information, instead taking shortcuts and relying on shallow, unsupported inferences; further, we design four tasks with different levels of difficulty to better delineate the models’ capability limits, along both the spatial dimension (from single- to multi-step spatial reasoning) and the temporal dimension (from keyframe- to sequence-baesd context), and observe a clear performance drop on both two dimensions. This reveals that both complex clue-based spatial reasoning and long-term memory retrieval are two distinct weaknesses that hinder the model’s performance on OST-Bench, highlighting the core challenges that must be addressed to advance online embodied reasoning.
@@ 2 Related Work
Spatial Reasoning Benchmarks. Early scene understanding benchmarksscanqa ; sqa3d ; sceneverse ; lyu2024mmscanmultimodal3dscene ; multi3drefer ; 3dvista introduced diverse task taxonomies to comprehensively evaluate various aspects of visual scene interpretation, with spatial understanding consistently recognized as the most fundamental component. Benchmarks such as ScanQAscanqa , SQA3Dsqa3d , SceneVersesceneverse , and MMScanlyu2024mmscanmultimodal3dscene emphasized semantic understanding and incorporated object locations and spatial relations, they largely treated spatial relationships as semantic attributes, focusing primarily on complex spatial semantics rather than explicitly targeting spatial reasoning. With the rapid advancement of Multimodal Large Language Models (MLLMs), recent benchmarks have begun to place greater emphasis on spatial reasoning evaluation, SpatialR-GPTspatialrgpt and CV-Benchcambrian1 require models to reason about 3D information, such as depth and distance from a single image, VSIyang2024thinkingspacemultimodallarge proposed a finer-grained categorization of spatial reasoning tasks, systematically evaluating models’ ability to infer 3D scene layouts from 2D video inputs, covering both relative and absolute spatial relationships. Existing spatial reasoning benchmarks predominantly operate in an offline setting, focusing on static scenes and requiring models to perform reasoning over a fixed set of images or videos of predefined length. In contrast, our OST-Bench adopts an online setting, emphasizing dynamic scene understanding from an agent-centric perspective, and offers an alternative perspective for evaluating spatial reasoning capabilities. It includes a wider range of complex question types to assess more diverse and fine-grained spatio-temporal reasoning abilities.
@@
MMScan lyu2024mmscanmultimodal3dscene Video/PC. Offline
@@ 3.1 Task Formulation
A static scene consists of a set of immobile objects, and understanding such a scene involves reasoning about individual entities and their relationshipslyu2024mmscanmultimodal3dscene , such as the object attribute(intrinsic properties of individual objects, including category, color, material, shape, size, and function), the attribute / spatial relationship between objects, and the spatial relationship between objects and a given viewpoint (provided either textually or via a virtual camera input). When a dynamic agent is introduced, it introduces new relational dynamics and opens up additional avenues for investigation. These can be categorized into three main categories that form the core focus of our benchmark evaluation: (1) Agent state: The position and orientation of the agent, which continuously change as the agent explores. (2) Agent visible info: The perceptual information available from the agent’s point of view at a given moment includes the existence of visible objects, their count, diversity, and the timing of their appearance. The information visible to the agent is continuously updated as the agent explores the scene. (3) Agent-object spatial relationship: 3D spatial relations between the agent and objects, described by relative or absolute distance/direction, constantly change as the agent explores.
@@ 3.2 Meta-dataset Collection and Processing
Base Dataset Acquisition. The three real scene datasets, ScanNetscannet , ARKitScenesarkitscenes and Matterport3Dmp3d , contain rich scene information along with RGB-D videos/images and their corresponding camera information, totaling 7.6k scenes. Building on this foundation, EmbodiedScanwang2023embodiedscan provided a large number of high-quality 9-DOF bounding box annotations for the objects in these scenes. MMScanlyu2024mmscanmultimodal3dscene further enriched these scenes with a large number of highly quality, manually annotated object- and region-level semantic annotations. We selected a total of 1.4k scenes from the validation/test splits of these three datasets and constructed our dataset based on the annotations from EmbodiedScan and MMScan.
@@ 4.1 Benchmark Models & Evaluation Metrics
We evaluate the performance of multiple multi-modal large language models (MLLMs), including both proprietary models (Claude-3.5-SonnetClaude3 , GPT-4ogpt4o , GPT-4.1gpt4 , Gemini-2.0-Flashgemini20 , and its thinking variant) and open-source models (InternVL-2.5internvl25 , QwenVL-2.5qwenvl25 , LLaVA-Onevisionllavaonevision , and LLaVA-Videollavavideo of different scales). Each model is tested in a zero-shot setting and conducts inference in a multi-turn dialogue format. To establish performance boundaries, we include two baselines: a human baseline and a chance-level baseline. For the human baseline, we ensure that participants have no prior exposure to the test scenes. The chance-level method adopts a random selection approach, randomly picks one answer from all possible choices for Judgement/Counting/Temporal-Localization questions, and for Estimation questions, it always outputs the mean value calculated from all potential numeric. As for the evaluation metrics, Judgement questions are considered correct if the model selects the same option as the ground truth. For Counting and Temporal Localization questions, the model’s output—whether a number or a turn index—must exactly match the ground truth to be deemed correct. For Estimation questions, we adopt the Mean Relative Accuracy (MRA) metric from VSIyang2024thinkingspacemultimodallarge to score the similarity between the model’s floating-point output and the ground truth.
@@ 4.2 Main Results
We report the performance of various models on our benchmark. Tab. 2 presents the performance of each model in all different subtype. Tab.3 summarizes the models’ overall performance, including their overall average scores, average scores for each of the three main categories, and performance across different question formats.
Substantial Gap between MLLMs’ and Human’s Performance. Model accuracy lags significantly behind human performance, with consistent gaps across all question types as shown in Tab.2. According to Tab.3, even the most advanced models lag behind human performance by over 30% in the overall average score. This performance gap remains substantial across three main task categories and all question formats. These findings suggest that current MLLMs fall short on OST-Bench, illustrating how our benchmark presents a novel challenge, demanding stronger online spatio-temporal perception and reasoning capabilities. This observation motivates us to investigate further the reasons for the subpar performance of the models on this benchmark.
@@ A.3 Rule-based Generation
  • Orientation (Judgement): This binary-choice question asks whether the agent has rotated clockwise or counterclockwise by an angle(less than 180 degrees) relative to its orientation at the end of round $T_{1}$. We compute the angle between the current orientation vector and the one at the end of $T_{1}$. To exclude ambiguous borderline cases, questions are generated only if the angle lies within the intervals $[\theta,180-\theta]$ or $[180+\theta,360-\theta]$, where $\theta$ is a threshold used to exclude borderline cases. Angles within the first interval indicate clockwise rotation, while those within the second indicate counterclockwise rotation.
@@ Appendix C Experiment Analysis Details
C.1 Cases of Three Error Types

### v2

@@ Abstract
Recent advances in multimodal large language models (MLLMs) have shown remarkable capabilities in integrating vision and language for complex reasoning. While most existing benchmarks evaluate models under offline settings with a fixed set of pre-recorded inputs, we introduce OST-Bench, a benchmark designed to evaluate Online Spatio-Temporal understanding from the perspective of an agent actively exploring a scene. The “Online” aspect emphasizes the need to process and reason over incrementally acquired observations, while the “Spatio-Temporal” component requires integrating current visual inputs with historical memory to support dynamic spatial reasoning. OST-Bench better reflects the challenges of real-world embodied perception. Built on an efficient data collection pipeline, OST-Bench consists of 1.4k scenes and 10k question-answer pairs collected from ScanNet, Matterport3D, and ARKitScenes. We evaluate several leading MLLMs on OST-Bench and observe that they fall short on tasks requiring complex spatio-temporal reasoning. Under the online setting, their accuracy declines as the exploration horizon extends and the memory grows. Through further experimental analysis, we identify common error patterns across models and find that both complex clue-based spatial reasoning demands and long-term memory retrieval requirements significantly drop model performance along two separate axes, highlighting the core challenges that must be addressed to improve online embodied reasoning. To foster further research and development in the field, our codes, dataset, and benchmark are available at https://github.com/InternRobotics/OST-Bench.
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs)bai2023qwentechnicalreport ; chen2024internvlscalingvisionfoundation ; liu2024improvedbaselinesvisualinstruction ; touvron2023llamaopenefficientfoundation ; lin2024vilapretrainingvisuallanguage ; zhang2024longcontexttransferlanguage ; li2024llavanextinterleavetacklingmultiimagevideo have shown remarkable capabilities in integrating vision and language for complex reasoning. However, most existing benchmarks scanrefer ; lyu2025mmscanmultimodal3dscene ; sqa3d ; scanqa ; sceneverse ; referit3d , evaluate models under offline settings, where reasoning is performed over a fixed set of pre-recorded inputs, such as reconstructed 3D scenes or images and videos, this does not capture the online nature of embodied tasks.
@@
To address this gap, we introduce OST-Bench, a benchmark designed to evaluate Online Spatio-Temporal understanding from the perspective of an agent actively exploring a scene. The term Online emphasizes the agent’s need to perceive, remember, and reason over incrementally received observations, rather than complete, pre-recorded scene data. The term Spatio-Temporal highlights the need to integrate current visual observation with historical memory to support dynamic spatial reasoning. To more accurately simulates real-world embodied perception, OST-Bench defines tasks from these perspectives: the agent, its surrounding environment, and their relationship, contains three main categories(Fig.2): (1) Agent State: the agent’s understanding of its own state, (2) Agent Visible Info: the agent’s dynamic interpretation of visible scene information, and (3)Agent-object Spatial Relationship: the agent’s dynamic understanding of spatial relationships with objects, all posed in an online, temporally grounded fashion. OST-Bench comprises 1.4k real-world scenes sourced from the test and validation splits of ScanNetscannet , Matterport3Dmp3d , and ARKitScenesarkitscenes , accompanied by 10k QA pairs covering a diverse range of subtypes. Our benchmark provides a rigorous testbed for assessing the online spatio-temporal reasoning ability of MLLMs in realistic, embodied settings.
@@
We evaluate leading MLLMs on OST-Bench and find its online spatio-temporal nature poses significant challenges for the models, even the most advanced models lag behind human performance by over 30%. Models perform poorly on tasks requiring complex spatio-temporal reasoning, with accuracy declining as exploration steps increase and memory grows under the online setting. Based on an in-depth experimental analysis, we observe a phenomenon which we term Spatio-temporal Reasoning Shortcut-when reasoning over long-term memory, models tend to avoid retrieving key information, instead taking shortcuts and relying on shallow, unsupported inferences; further, we design four tasks with different levels of difficulty to better delineate the models’ capability limits, along both the spatial dimension (from single- to multi-step spatial reasoning) and the temporal dimension (from keyframe- to sequence-baesd context), and observe a clear performance drop on both dimensions. This reveals that both complex clue-based spatial reasoning and long-term memory retrieval are two distinct weaknesses that hinder the model’s performance on OST-Bench, highlighting the core challenges that must be addressed to advance online embodied reasoning. Moreover, our fine-tuning analysis shows that data-driven training alone yields only limited improvement, suggesting that further progress will likely require advances in model architecture and training methodology rather than sheer data scaling.
@@ 2 Related Work
Spatial Reasoning Benchmarks. Early scene understanding benchmarksscanqa ; sqa3d ; sceneverse ; lyu2025mmscanmultimodal3dscene ; multi3drefer ; 3dvista introduced diverse task taxonomies to comprehensively evaluate various aspects of visual scene interpretation, with spatial understanding consistently recognized as the most fundamental component. Benchmarks such as ScanQAscanqa , SQA3Dsqa3d , SceneVersesceneverse , and MMScanlyu2025mmscanmultimodal3dscene emphasized semantic understanding and incorporated object locations and spatial relations, they largely treated spatial relationships as semantic attributes, focusing primarily on complex spatial semantics rather than explicitly targeting spatial reasoning. With the rapid advancement of Multimodal Large Language Models (MLLMs), recent benchmarks have begun to place greater emphasis on spatial reasoning evaluation, SpatialR-GPTspatialrgpt and CV-Benchcambrian1 require models to reason about 3D information, such as depth and distance from a single image, VSIyang2024thinkingspacemultimodallarge proposed a finer-grained categorization of spatial reasoning tasks, systematically evaluating models’ ability to infer 3D scene layouts from 2D video inputs, covering both relative and absolute spatial relationships. Existing spatial reasoning benchmarks predominantly operate in an offline setting, focusing on static scenes and requiring models to perform reasoning over a fixed set of images or videos of predefined length. In contrast, our OST-Bench adopts an online setting, emphasizing dynamic scene understanding from an agent-centric perspective, and offers an alternative perspective for evaluating spatial reasoning capabilities. It includes a wider range of complex question types to assess more diverse and fine-grained spatio-temporal reasoning abilities.
@@
MMScan lyu2025mmscanmultimodal3dscene Video/PC. Offline
@@ 3.1 Task Formulation
A static scene consists of a set of immobile objects, and understanding such a scene involves reasoning about individual entities and their relationshipslyu2025mmscanmultimodal3dscene , such as the object attribute(intrinsic properties of individual objects, including category, color, material, shape, size, and function), the attribute / spatial relationship between objects, and the spatial relationship between objects and a given viewpoint (provided either textually or via a virtual camera input). When a dynamic agent is introduced, it introduces new relational dynamics and opens up additional avenues for investigation. These can be categorized into three main categories that form the core focus of our benchmark evaluation: (1) Agent state: The position and orientation of the agent, which continuously change as the agent explores. (2) Agent visible info: The perceptual information available from the agent’s point of view at a given moment includes the existence of visible objects, their count, diversity, and the timing of their appearance. The information visible to the agent is continuously updated as the agent explores the scene. (3) Agent-object spatial relationship: 3D spatial relations between the agent and objects, described by relative or absolute distance/direction, constantly change as the agent explores.
@@ 3.2 Meta-dataset Collection and Processing
Base Dataset Acquisition. The three real scene datasets, ScanNetscannet , ARKitScenesarkitscenes and Matterport3Dmp3d , contain rich scene information along with RGB-D videos/images and their corresponding camera information, totaling 7.6k scenes. Building on this foundation, EmbodiedScanwang2023embodiedscan provided a large number of high-quality 9-DOF bounding box annotations for the objects in these scenes. MMScanlyu2025mmscanmultimodal3dscene further enriched these scenes with a large number of highly quality, manually annotated object- and region-level semantic annotations. We selected a total of 1.4k scenes from the validation/test splits of these three datasets and constructed our dataset based on the annotations from EmbodiedScan and MMScan.
@@ 4.1 Benchmark Models & Evaluation Metrics
We evaluate the performance of multiple multi-modal large language models (MLLMs), including both proprietary models (Claude-3.5-SonnetClaude3 , GPT-4ogpt4o , GPT-4.1gpt4 , Gemini-2.0-Flashgemini20 , and its thinking variant) and open-source models (InternVL-2.5internvl25 , QwenVL-2.5qwenvl25 , LLaVA-Onevisionllavaonevision , and LLaVA-Videollavavideo of different scales). Each model is tested in a zero-shot setting and conducts inference in a multi-turn dialogue format. (In addition to these general-purpose VLMs, we also include several models specifically designed with explicit spatial grounding or memory mechanisms to varying degrees, their results are reported in Appendix C.1.) To establish performance boundaries, we include two baselines: a human baseline and a chance-level baseline. For the human baseline, we ensure that participants have no prior exposure to the test scenes. The chance-level method adopts a random selection approach, randomly picks one answer from all possible choices for Judgement/Counting/Temporal-Localization questions, and for Estimation questions, it always outputs the mean value calculated from all potential numeric. As for the evaluation metrics, Judgement questions are considered correct if the model selects the same option as the ground truth. For Counting and Temporal Localization questions, the model’s output—whether a number or a turn index—must exactly match the ground truth to be deemed correct. For Estimation questions, we adopt the Mean Relative Accuracy (MRA) metric from VSIyang2024thinkingspacemultimodallarge to score the similarity between the model’s floating-point output and the ground truth.
@@ 4.2 Main Results
We report the performance of various models on our benchmark. Tab. 2 presents the performance of each model in all different subtype. Tab.3 summarizes the models’ overall performance, including their overall average scores, average scores for each of the three main categories, and performance across different question formats. Additional model results and further analysis are provided in Appendix C.1.
Substantial Gap between MLLMs’ and Human’s Performance. Model accuracy lags significantly behind human performance, with consistent gaps across all question types as shown in Tab.2. According to Tab.3, even the most advanced models lag behind human performance by nearly 30% in the overall average score. This performance gap remains substantial across three main task categories and all question formats. These findings suggest that current MLLMs fall short on OST-Bench, illustrating how our benchmark presents a novel challenge, demanding stronger online spatio-temporal perception and reasoning capabilities. This observation motivates us to investigate further the reasons for the subpar performance of the models on this benchmark.
@@ 4.3.2 Cross-View Analysis
4.3.3 Fine-tuning Analysis
To dig deeper into the upper bound of current models’ capabilities and to better understand how much of the performance gap can be recovered through training with in-domain data, we conducted fine-tuning experiments on several representative models. Following a procedure similar to that used for constructing the benchmark samples, we generated training data from 7k training scenes across ScanNet, Matterport3D, and ARKitScenes, yielding a total of 50k annotated samples. All models were fine-tuned for a single epoch, and the results are summarized in Tab. 4. Overall, all evaluated models achieved performance gains exceeding 10% after fine-tuning. However, a closer analysis reveals several important insights:
Method Setting Overall JUD. EST. CNT. Temp-Loc. A State A Info AO
Zero-Shot 41.2 50.1 22.7 62.1 36.6 40.4 52.6 31.5
QwenVL2.5-7B Fine-Tuned 54.0 59.0 41.2 74.6 50.2 48.3 69.8 43.5
Zero-Shot 44.6 52.4 29.3 56.4 49.2 41.9 58.6 34.6
InternVL2.5-8B Fine-Tuned 57.4 64.1 38.5 74.9 57.5 44.0 79.3 46.3
Zero-Shot 50.8 61.4 31.5 61.1 45.9 45.4 73.9 34.0
InternVL2.5-38B Fine-Tuned 60.2 68.4 44.1 73.1 56.1 50.8 81.7 47.5
Table 4: Performance comparison of models under zero-shot and fine-tuned settings. "A." abbrev for "Agent" and "AO." abbrev for "Agent-Object Spatial Relationship".
  • Among the three major task categories, the largest improvements emerged in Agent Visible Info tasks — particularly for models with smaller parameter sizes. In contrast, the other two task categories, despite showing some gains, remained at or below 50% accuracy. This indicates that even with in-domain adaptation, models still struggle with tasks that demand complex spatio-temporal reasoning. Simple supervised fine-tuning on OST-Bench is therefore insufficient to resolve its core challenges.
  • All four question formats benefited from fine-tuning, yet deeper inspection of predictions reveals more nuanced observations. Although fine-tuning improves the scores of Estimation (EST) and Judgement (JUD) tasks, closer inspection reveals that these gains do not reflect genuine reasoning improvements: models frequently output nearly identical values or default to the same option across samples, indicating reliance on dataset-specific shortcuts or memorization rather than true understanding. Moreover, their instruction-following ability degrades post-finetuning, with many responses failing to provide both the final answer and the required reasoning.
While fine-tuning significantly improves raw performance, a considerable gap remains compared to human-level accuracy. This highlights two key points: (1) data-only supervised fine-tuning is insufficient to solve the challenges posed by OST-Bench — improvements may also be required on the model architecture or training methodology side; and (2) the benchmark itself is both challenging and robust. Despite being constructed using templates, OST-Bench resists shortcut learning and cannot be easily exploited through superficial patterns in the training distribution.
@@ 5 Limitations and Conclusion
6 Acknowledgement
This work is funded in part by the National Key R&D Program of China, and Shanghai Artificial Intelligence Laboratory.
@@ A.3 Rule-based Generation
  • Orientation (Judgement): This binary-choice question asks whether the agent has rotated clockwise or counterclockwise by an angle(less than 180 degrees) relative to its orientation at the end of round $T_{1}$. We compute the angle between the current orientation vector and the one at the end of $T_{1}$. To exclude ambiguous borderline cases, questions are generated only if the angle lies within the intervals $\theta,180-\theta]$ or $180+\theta,360-\theta]$, where $\theta$ is a threshold used to exclude borderline cases. Angles within the first interval indicate clockwise rotation, while those within the second indicate counterclockwise rotation.
@@ Appendix C Experiment Analysis Details
C.1 Spatially-Grounded Model Evaluation
In addition to general-purpose VLMs, we further evaluate several representative models that incorporate spatial grounding and memory mechanisms to varying degrees, including Spatial-MLLMwu2025spatialmllmboostingmllmcapabilities , VLM-3Rfan2025vlm3rvisionlanguagemodelsaugmented , and LLaVA-3Dzhu2024llava3dsimpleeffectivepathway .
  • Spatial-MLLM and VLM-3R follow a VGGTwang2025vggtvisualgeometrygrounded + VLM architecture and take RGB image sequences as input, where VGGT provides geometry-aware scene representations.
  • LLaVA-3D leverages RGB-D image sequences to encode 3D spatial information into 2D token embeddings.
All three models were trained on spatial reasoning datasets and achieved strong results on their respective benchmarks. We evaluate their performance on OST-Bench and compare them to their corresponding base models — Spatial-MLLM vs. QwenVL2.5-3B, and VLM-3R / LLaVA-3D vs. LLaVA-Video-7B.(Tab 5) The key findings are summarized below:
  • Only VLM-3R delivers consistent gains over its base model. Spatial-MLLM and LLaVA-3D exhibit substantial performance drops, whereas VLM-3R shows steady improvements, particularly in Agent State, Agent-Object Spatial Relationship, and Estimation tasks.
  • Instruction-following ability degrades noticeably. Compared to their base models—which reliably follow prompts and output both answers and reasoning—all three grounded models struggle to adhere to the required response format. Spatial-MLLM is constrained to producing only floating-point values or multiple-choice options, while all three models frequently omit reasoning or generate incoherent explanations.
  • Limited generalization beyond training-aligned distributions. Despite excelling on spatial reasoning datasets such as VSI and MMScan, Spatial-MLLM and LLaVA-3D fail to generalize effectively to OST-Bench, which features more diverse and temporally grounded prompts. VLM-3R demonstrates partial transferability, yet its gains remain modest.
These observations suggest that while memory-enhanced spatial grounding can improve performance on tasks aligned with model pretraining objectives, it often comes at the expense of generalization. Such models may lose part of the base LLM’s robustness — underperforming on previously simple tasks (e.g., Agent Visible Info or Counting questions in OST-Bench) and struggling with instruction-following in out-of-distribution settings. They tend to excel only on in-domain tasks and transfer poorly to broader benchmarks like OST-Bench, which involve more diverse and complex reasoning demands.
Method Overall JUD. EST. CNT. TEMP. A. State A. Info AO.
(base) QwenVL2.5-3B 34.8 47.9 18.7 59.4 19.8 34.2 47.5 25.7
Spatial-MLLM 26.8 37.3 21.9 29.5 15.3 25.5 39.4 20.9
(base) LLaVA-Video-7B 39.3 52.8 16.1 63.1 33.8 33.5 58.3 28.8
VLM-3R 42.9 55.1 28.3 49.6 36.0 39.9 58.1 34.4
LLaVA-3D 30.1 46.1 5.9 13.5 36.3 29.7 38.4 26.3
Table 5: Performance comparison between specially designed models and their corresponding base models. "A." abbrev for "Agent" and "AO." abbrev for "Agent-Object Spatial Relationship".
C.2 More Findings in Tables
C.3 Cases of Three Error Types